OpenStudy (anonymous):

For every metre below the surface of a fluid, the light intensity is reduced by 5%. At what depth is the light intensity 25% of the light intensity at the surface? (accurate to the nearest metre)

OpenStudy (anonymous):

\[1metre=0.05\] \[?metre=1-0.25\] \[?metre=0.75\] Cross multiply with the first line and third. \[1\times0.75=0.75\] \[\frac{ 0.75 }{ 0.05 }=15\] Therefore 15 metres. Not sure if it's correct but if you have the answers, please do tell me if it's right or wrong.

OpenStudy (kropot72):

\[I _{d}=I _{s}(1-0.05)^{d}\] \[I _{d}=intensity\ at\ d\ metres\] \[I _{s}=intensity\ at\ surface\] \[d=depth\ (metres)\] \[\frac{I _{d}}{I _{s}}=0.25=0.95^{d}\] Taking natural logs of both sides gives: \[\ln 0.25=d \times \ln 0.95\] \[d=\frac{\ln 0.25}{\ln 0.95}=you\ can\ calculate\]